# Scaling over Scaling: Exploring Test-Time Scaling Pareto in Large Reasoning Models

### v1

@@ Abstract
Large reasoning models (LRMs) have exhibited the capacity of enhancing reasoning performance via internal test-time scaling. Building upon this, a promising direction is to further scale test-time compute to unlock even greater reasoning capabilities. However, as we push these scaling boundaries, systematically understanding the practical limits and achieving optimal resource allocation becomes a critical challenge. In this paper, we investigate the scaling Pareto of test-time scaling and introduce the Test-Time Scaling Performance Model (TTSPM). We theoretically analyze two fundamental paradigms for such extended scaling, parallel scaling and sequential scaling, from a probabilistic modeling perspective. Our primary contribution is the derivation of the saturation point on the scaling budget for both strategies, identifying thresholds beyond which additional computation yields diminishing returns. Remarkably, despite their distinct mechanisms, both paradigms converge to a unified mathematical structure in their upper bounds. We empirically validate our theoretical findings on challenging reasoning benchmarks, including AIME, MATH-500, and GPQA, demonstrating the practical utility of these bounds for test-time resource allocation. We hope that this work provides insights into the cost-benefit trade-offs of test-time scaling, guiding the development of more resource-efficient inference strategies for large reasoning models.
@@ 1 Introduction
Building upon the foundation of internally scaled reasoning, we naturally pose the next question: Can these already powerful LRMs be further enhanced by allocating even more compute during inference? This inquiry motivates our exploration of test-time scaling over internally scaled LRMs, a concept we refer to as scaling over scaling. However, as we consider increasing test-time compute, such as by generating additional solution candidates or performing more rounds of iterative refinement, it becomes essential to ask: Is such scaling always beneficial? In other words, does a point of diminishing returns exist, beyond which additional computational effort yields negligible performance gains? Identifying and characterizing this scaling Pareto is critical for preventing resource over-allocation, especially in most TTS strategies where the absence of supervised feedback complicates efficient compute usage.
@@
To address the above questions, we introduce the Test-Time Scaling Performance Model (TTSPM), a theoretical probabilistic model that reimagines test-time scaling performance through the lens of probabilistic modeling (as detailed in Section 3). TTSPM begins with two fundamental TTS paradigms: 1) parallel scaling (wang2023self, ), where multiple reasoning paths and solutions are generated independently; and 2) sequential scaling (tian2025think, ), where a solution is iteratively refined round by round. Their performance gains with increasing compute often exhibit a characteristic saturation curve. We develop a general probabilistic model that captures this behavior, positing that the probability of success with the scaling budget $N$ approaches a maximum achievable performance $F_{\text{max}}$ as $N$ grows. Based on this probabilistic model, we analyze the marginal performance gain per additional generation, which is referred to as a scaling unit. We then formally define the scaling Pareto as the point where this marginal gain is no longer higher than a predefined small threshold $\epsilon$ as increasing scaling units. Using this definition, we derive a unified saturation point for the scaling budget, which signifies the critical point beyond which further scaling is deemed inefficient.
@@
We empirically validate the practical utility of our theoretical bounds on challenging mathematical reasoning and PhD-level science question answering benchmarks, including AIME (maa2024aime, ), MATH-500 (lightman2024let, ), and GPQA (rein2024gpqa, ). Our experiments (see Section 4) demonstrate that the derived bounds effectively predict the onset of the scaling Pareto, aligning with observed performance saturation. These findings offer valuable insights. First, LRMs can indeed be further improved via simple, verifier-free test-time scaling. Second, this improvement is not limitless and exhibits a predictable saturation point. Our TTSPM provides a principled method to estimate this point, enabling more efficient resource allocation by avoiding unnecessary token consumption.
@@
Overall, the contributions of this paper are as follows: 1) We propose TTSPM, a theoretical probabilistic model that effectively characterizes test-time scaling Pareto in large reasoning models. 2) With TTSPM, we derive a general upper bound for scaling units. It is applicable to both parallel and sequential scaling strategies, revealing a consistent mathematical structure underlying their saturation behavior. 3) We outline how this theoretical bound can be empirically validated and discuss its practical implications for optimizing test-time compute, thereby guiding the development of more resource-efficient inference scaling strategies for large reasoning models.
@@ 2 Related Work
However, a fundamental question that remains underexplored is the inherent limit of performance gain from these test-time scaling strategies, particularly for verifier-free approaches. While it is intuitive that more computation should lead to better performance, the relationship is unlikely to be linear indefinitely. Understanding when these scaling efforts hit a point of diminishing returns—a scaling Pareto—is critical for efficient resource utilization. Our work attempts to address this gap by theoretically modeling this trend and deriving explicit upper bounds (saturation points) on the computational budget for both parallel and sequential scaling, beyond which further scaling offers minimal performance gains. This provides a principled basis for optimizing test-time compute allocation, complementing existing empirical explorations of TTS efficiency.
@@ 3.1 Probabilistic Modeling of Test-Time Scaling
3.2 Performance Function to Scaling Pareto
@@
Derivation of Scaling Pareto.
We define the scaling Pareto as the operational region where this marginal performance gain, $\Delta F(N)$, becomes practically insignificant. Specifically, we posit that further scaling is inefficient if $\Delta F(N)$ falls below a predefined small positive threshold $\epsilon$. This threshold $\epsilon$ represents the minimum performance improvement deemed worthwhile for the cost of an additional scaling unit. Our objective is to identify an upper bound for the scaling budget, $N_{\text{upper}}$, such that for any $N\geq N_{\text{upper}}$, the marginal gain $\Delta F(N)$ is consistently less than the chosen threshold $\epsilon$. This condition is formally expressed as:
@@
This $N^{*}$ signifies the scaling budget beyond which scaling additional units is expected to yield performance improvements below the threshold $\epsilon$, effectively marking the scaling Pareto.
Insights into the Scaling Pareto.
The core finding of this paper is that both parallel and sequential test-time scaling strategies, despite their operational differences, exhibit a performance saturation phenomenon—the scaling Pareto—that can be characterized by a unified mathematical structure, as induced in Eq. (14). This suggests a fundamental principle at play: as more computational units ($N$) are invested at test time, the probability of encountering an unsolved problem that can be successfully addressed by the next unit diminishes. The exponential decay term $(1-p_{x})^{N}$ in our marginal gain formula (Eq. (9)) captures this essence. The derived saturation point, $N^{*}$, provides a quantitative tool to identify when the expected marginal gain drops below a practical threshold $\epsilon$, signaling that further scaling is unlikely to be cost-effective.
@@ 4.2 Main Results of Test-Time Scaling
4.3 Analysis of the Scaling Pareto
Figure 2 shows the scaling curves across different benchmarks, model sizes, and scaling strategies, exhibiting a consistent pattern of initial rapid improvement followed by a Pareto, confirming our theoretical model. We conclude that several important insights emerge from these curves: 1) The scaling Pareto is a universal phenomenon observed across all experimental settings. For both the 1.5B and 7B models, performance improvements become increasingly marginal as the number of generations increases, eventually reaching a point where additional scaling yields negligible gains. This empirical observation strongly validates our theoretical derivation of the scaling Pareto. 2) Through test-time scaling, a smaller model can approach and sometimes even surpass the performance of a larger model without scaling. On MATH-500, the 1.5B model with parallel scaling at $N=32$ achieves 83.6% accuracy, exceeding the 7B model’s vanilla performance of 82.0%. This demonstrates that test-time scaling can be a cost-effective alternative to model size scaling, offering a practical approach to enhance reasoning capabilities without the computational burden of training larger models. 3) The distribution discrepancy of different benchmark results in distinct scaling characteristics. GPQA shows the steepest initial improvement, particularly for parallel scaling, suggesting that this benchmark benefits most from diverse solution exploration. In contrast, AIME2025 exhibits more gradual improvements, indicating that the problems in this benchmark may require more focused refinement rather than broad exploration. Further accuracy scaling curves presented in Appendix C demonstrate similar insights.
@@ 4.3 Analysis of the Scaling Pareto
4.4 Verification of the Scaling Pareto
To validate our theoretical model, TTSPM, we compare the predicted scaling Pareto with the empirically observed optimal scaling budget for each problem in the MATH-500 dataset. To achieve a theoretical prediction of the scaling Pareto of $N$, we should first estimate the value of the probability $p_{x}$, so that we can employ Eq. (14) to calculate the saturation point of the scaled units. We randomly split the MATH-500 dataset into a “validation set” and a “test set” in a proportion of 8:2. We then utilize the “validation set” to estimate the unknown entry $\frac{\epsilon}{F_{\max}}$. With this, we can calculate the predicted scaling Pareto on the test set. Further details are presented in Appendix A.
Figure LABEL:fig:verification presents the correlation between the theoretically predicted and the practically observed scaling saturation points in the MATH-500 dataset. The strong positive correlations observed in all different models and scaling strategies provide compelling evidence for the predictive power of our theoretical model. Parallel scaling shows particularly strong correlations, with Pearson’s correlation coefficient $r$ = 0.743 for the 1.5B model and $r$ = 0.802 for the 7B model. These correlations indicate that our theoretical model can effectively predict the point at which additional scaling becomes inefficient, allowing us to make informed decisions about computational resource allocation. The stronger correlation observed for parallel scaling suggests that our model is particularly well-suited for predicting the scaling behavior of independent sampling approaches. Interestingly, we observe that the 7B model with parallel scaling (Figure LABEL:fig:scatter_plot-7b_para) shows the highest correlation, indicating that larger, more stable models may be more predictable in their scaling behavior. This finding has important implications for resource allocation in large-scale deployments, where accurate prediction of scaling benefits can lead to significant computational savings.
@@ 5 Conclusion and Future Work
In this paper, we investigate the scaling Pareto phenomenon in test-time computation for large reasoning models, developing a unified theoretical test-time scaling performance model (TTSPM) that characterizes the point at which additional computational units yield diminishing returns. Our key contribution is the derivation of a general upper bound (saturation point) with identical mathematical structure for both parallel and sequential scaling strategies, suggesting a fundamental principle governs test-time scaling limits regardless of mechanism. Experiments on challenging reasoning benchmarks validate our theoretical bounds, enabling practitioners to make informed decisions about cost-benefit trade-offs in test-time scaling. In the future, we aim to enhance our method in several key aspects. Currently, our TTSPM is task- and model-specific, requiring careful instantiation and parameter estimation for each new setting. To address this limitation, we plan to adapt TTSPM to support rapid parameter estimation across diverse scenarios. Additionally, we intend to explore hybrid scaling strategies, develop dynamic parameter estimation techniques, and extend the core principles of our approach to a broader range of reasoning tasks.
@@ Appendix A Experimental Details
These estimations are performed for every problem in both the validation and test sets, providing the problem-specific parameters needed for our scaling Pareto predictions.
@@ B.2 Detailed Derivation of the Inequality for N
We start from the condition for the scaling Pareto (Eq ( 10)):
@@
For this inequality to be meaningful in the context of finding a positive $N$ where performance gain diminishes, we require the right-hand side to be less than 1 (since $(1-p_{x})^{N}$ will be less than 1 for $N\geq 1$ and $10<p_{x}<$ 1). If $\frac{\epsilon}{F_{\max}\cdot p_{x}}\geq 1$, it means that even for $N=0$ (or $N=1$, depending on interpretation), the marginal gain $F_{\max}\cdot p_{x}$ is already less than or equal to $\epsilon$, implying the Pareto is reached immediately. Thus, we assume $10<\frac{\epsilon}{F_{\max}\cdot p_{x}}<$ 1, which implies $\epsilon<F_{\max}\cdot p_{x}$.
@@ Appendix C Additional Experimental Results
Figure LABEL:fig:acc_curve presents accuracy scaling curves across four benchmarks under different scaling strategies (parallel vs. sequential) and model sizes (1.5B vs. 7B). Across all settings, we observe a consistent trend: initial improvements in accuracy with increased generations per problem ($N$), followed by diminishing returns, forming a clear empirical scaling Pareto frontier.
In addition to the fact that the scaling Pareto phenomenon is consistent and universal across tasks, models, and strategies, the above results further lead to several key insights: 1) Test-time scaling enables smaller models to rival or outperform larger, unscaled models. For example, the 1.5B model with sufficient scaling often approaches or exceeds the performance of the 7B vanilla baseline, particularly under parallel scaling. This highlights the practical value of test-time scaling as a lightweight alternative to model scale. 2) Parallel scaling consistently outperforms sequential scaling in terms of final accuracy. While both strategies benefit from increased $N$, parallel scaling achieves better asymptotic performance and exhibits more stable behavior across all benchmarks. These findings not only align with our theoretical predictions but also suggest that scaling compute at test time, especially via parallel strategies, can be a powerful tool for improving performance without increasing model size.

### v2

# Scaling over Scaling: Exploring Test-Time Scaling Plateau in Large Reasoning Models

@@ Abstract
Large reasoning models (LRMs) have exhibited the capacity of enhancing reasoning performance via internal test-time scaling. Building upon this, a promising direction is to further scale test-time compute to unlock even greater reasoning capabilities. However, as we push these scaling boundaries, systematically understanding the practical limits and achieving optimal resource allocation becomes a critical challenge. In this paper, we investigate the scaling plateau of test-time scaling and introduce the Test-Time Scaling Performance Model (TTSPM). We theoretically analyze two fundamental paradigms for such extended scaling, parallel scaling and sequential scaling, from a probabilistic modeling perspective. Our primary contribution is the derivation of the saturation point on the scaling budget for both strategies, identifying thresholds beyond which additional computation yields diminishing returns. Remarkably, despite their distinct mechanisms, both paradigms converge to a unified mathematical structure in their upper bounds. We empirically validate our theoretical findings on challenging reasoning benchmarks, including AIME, MATH-500, and GPQA, demonstrating the practical utility of these bounds for test-time resource allocation. We hope that this work provides insights into the cost-benefit trade-offs of test-time scaling, guiding the development of more resource-efficient inference strategies for large reasoning models.
@@ 1 Introduction
Building upon the foundation of internally scaled reasoning, we naturally pose the next question: Can these already powerful LRMs be further enhanced by allocating even more compute during inference? This inquiry motivates our exploration of test-time scaling over internally scaled LRMs, a concept we refer to as scaling over scaling. However, as we consider increasing test-time compute, such as by generating additional solution candidates or performing more rounds of iterative refinement, it becomes essential to ask: Is such scaling always beneficial? In other words, does a point of diminishing returns exist, beyond which additional computational effort yields negligible performance gains? Identifying and characterizing this scaling plateau is critical for preventing resource over-allocation, especially in most TTS strategies where the absence of supervised feedback complicates efficient compute usage.
@@
To address the above questions, we introduce the Test-Time Scaling Performance Model (TTSPM), a theoretical probabilistic model that reimagines test-time scaling performance through the lens of probabilistic modeling (as detailed in Section 3). TTSPM begins with two fundamental TTS paradigms: 1) parallel scaling (wang2023self, ), where multiple reasoning paths and solutions are generated independently; and 2) sequential scaling (tian2025think, ), where a solution is iteratively refined round by round. Their performance gains with increasing compute often exhibit a characteristic saturation curve. We develop a general probabilistic model that captures this behavior, positing that the probability of success with the scaling budget $N$ approaches a maximum achievable performance $F_{\text{max}}$ as $N$ grows. Based on this probabilistic model, we analyze the marginal performance gain per additional generation, which is referred to as a scaling unit. We then formally define the scaling plateau as the point where this marginal gain is no longer higher than a predefined small threshold $\epsilon$ as increasing scaling units. Using this definition, we derive a unified saturation point for the scaling budget, which signifies the critical point beyond which further scaling is deemed inefficient.
@@
We empirically validate the practical utility of our theoretical bounds on challenging mathematical reasoning and PhD-level science question answering benchmarks, including AIME (maa2024aime, ), MATH-500 (lightman2024let, ), and GPQA (rein2024gpqa, ). Our experiments (see Section 4) demonstrate that the derived bounds effectively predict the onset of the scaling plateau, aligning with observed performance saturation. These findings offer valuable insights. First, LRMs can indeed be further improved via simple, verifier-free test-time scaling. Second, this improvement is not limitless and exhibits a predictable saturation point. Our TTSPM provides a principled method to estimate this point, enabling more efficient resource allocation by avoiding unnecessary token consumption.
@@
Overall, the contributions of this paper are as follows: 1) We propose TTSPM, a theoretical probabilistic model that effectively characterizes test-time scaling plateau in large reasoning models. 2) With TTSPM, we derive a general upper bound for scaling units. It is applicable to both parallel and sequential scaling strategies, revealing a consistent mathematical structure underlying their saturation behavior. 3) We outline how this theoretical bound can be empirically validated and discuss its practical implications for optimizing test-time compute, thereby guiding the development of more resource-efficient inference scaling strategies for large reasoning models.
@@ 2 Related Work
However, a fundamental question that remains underexplored is the inherent limit of performance gain from these test-time scaling strategies, particularly for verifier-free approaches. While it is intuitive that more computation should lead to better performance, the relationship is unlikely to be linear indefinitely. Understanding when these scaling efforts hit a point of diminishing returns—a scaling plateau—is critical for efficient resource utilization. Our work attempts to address this gap by theoretically modeling this trend and deriving explicit upper bounds (saturation points) on the computational budget for both parallel and sequential scaling, beyond which further scaling offers minimal performance gains. This provides a principled basis for optimizing test-time compute allocation, complementing existing empirical explorations of TTS efficiency.
@@ 3.1 Probabilistic Modeling of Test-Time Scaling
3.2 Performance Function to Scaling Plateau
@@
Derivation of Scaling Plateau.
We define the scaling plateau as the operational region where this marginal performance gain, $\Delta F(N)$, becomes practically insignificant. Specifically, we posit that further scaling is inefficient if $\Delta F(N)$ falls below a predefined small positive threshold $\epsilon$. This threshold $\epsilon$ represents the minimum performance improvement deemed worthwhile for the cost of an additional scaling unit. Our objective is to identify an upper bound for the scaling budget, $N_{\text{upper}}$, such that for any $N\geq N_{\text{upper}}$, the marginal gain $\Delta F(N)$ is consistently less than the chosen threshold $\epsilon$. This condition is formally expressed as:
@@
This $N^{*}$ signifies the scaling budget beyond which scaling additional units is expected to yield performance improvements below the threshold $\epsilon$, effectively marking the scaling plateau.
Insights into the Scaling Plateau.
The core finding of this paper is that both parallel and sequential test-time scaling strategies, despite their operational differences, exhibit a performance saturation phenomenon—the scaling plateau—that can be characterized by a unified mathematical structure, as induced in Eq. (14). This suggests a fundamental principle at play: as more computational units ($N$) are invested at test time, the probability of encountering an unsolved problem that can be successfully addressed by the next unit diminishes. The exponential decay term $(1-p_{x})^{N}$ in our marginal gain formula (Eq. (9)) captures this essence. The derived saturation point, $N^{*}$, provides a quantitative tool to identify when the expected marginal gain drops below a practical threshold $\epsilon$, signaling that further scaling is unlikely to be cost-effective.
@@ 4.2 Main Results of Test-Time Scaling
4.3 Analysis of the Scaling Plateau
Figure 2 shows the scaling curves across different benchmarks, model sizes, and scaling strategies, exhibiting a consistent pattern of initial rapid improvement followed by a plateau, confirming our theoretical model. We conclude that several important insights emerge from these curves: 1) The scaling plateau is a universal phenomenon observed across all experimental settings. For both the 1.5B and 7B models, performance improvements become increasingly marginal as the number of generations increases, eventually reaching a point where additional scaling yields negligible gains. This empirical observation strongly validates our theoretical derivation of the scaling plateau. 2) Through test-time scaling, a smaller model can approach and sometimes even surpass the performance of a larger model without scaling. On MATH-500, the 1.5B model with parallel scaling at $N=32$ achieves 83.6% accuracy, exceeding the 7B model’s vanilla performance of 82.0%. This demonstrates that test-time scaling can be a cost-effective alternative to model size scaling, offering a practical approach to enhance reasoning capabilities without the computational burden of training larger models. 3) The distribution discrepancy of different benchmark results in distinct scaling characteristics. GPQA shows the steepest initial improvement, particularly for parallel scaling, suggesting that this benchmark benefits most from diverse solution exploration. In contrast, AIME2025 exhibits more gradual improvements, indicating that the problems in this benchmark may require more focused refinement rather than broad exploration. Further accuracy scaling curves presented in Appendix C demonstrate similar insights.
@@ 4.3 Analysis of the Scaling Plateau
4.4 Verification of the Scaling Plateau
To validate our theoretical model, TTSPM, we compare the predicted scaling plateau with the empirically observed optimal scaling budget for each problem in the MATH-500 dataset. To achieve a theoretical prediction of the scaling plateau of $N$, we should first estimate the value of the probability $p_{x}$, so that we can employ Eq. (14) to calculate the saturation point of the scaled units. We randomly split the MATH-500 dataset into a “validation set” and a “test set” in a proportion of 8:2. We then utilize the “validation set” to estimate the unknown entry $\frac{\epsilon}{F_{\max}}$. With this, we can calculate the predicted scaling plateau on the test set. Further details are presented in Appendix A.
Figure LABEL:fig:verification presents the correlation between the theoretically predicted and the practically observed scaling saturation points in the MATH-500 dataset. The strong positive correlations observed in all different models and scaling strategies provide compelling evidence for the predictive power of our theoretical model. Parallel scaling shows particularly strong correlations, with Pearson’s correlation coefficient $r$ = 0.768 for the 1.5B model and $r$ = 0.803 for the 7B model. These correlations indicate that our theoretical model can effectively predict the point at which additional scaling becomes inefficient, allowing us to make informed decisions about computational resource allocation. The stronger correlation observed for parallel scaling suggests that our model is particularly well-suited for predicting the scaling behavior of independent sampling approaches. Interestingly, we observe that the 7B model with parallel scaling (Figure LABEL:fig:scatter_plot-7b_para) shows the highest correlation, indicating that larger, more stable models may be more predictable in their scaling behavior. This finding has important implications for resource allocation in large-scale deployments, where accurate prediction of scaling benefits can lead to significant computational savings.
@@ 5 Conclusion and Future Work
In this paper, we investigate the scaling plateau phenomenon in test-time computation for large reasoning models, developing a unified theoretical test-time scaling performance model (TTSPM) that characterizes the point at which additional computational units yield diminishing returns. Our key contribution is the derivation of a general upper bound (saturation point) with identical mathematical structure for both parallel and sequential scaling strategies, suggesting a fundamental principle governs test-time scaling limits regardless of mechanism. Experiments on challenging reasoning benchmarks validate our theoretical bounds, enabling practitioners to make informed decisions about cost-benefit trade-offs in test-time scaling. In the future, we aim to enhance our method in several key aspects. Currently, our TTSPM is task- and model-specific, requiring careful instantiation and parameter estimation for each new setting. To address this limitation, we plan to adapt TTSPM to support rapid parameter estimation across diverse scenarios. Additionally, we intend to explore hybrid scaling strategies, develop dynamic parameter estimation techniques, and extend the core principles of our approach to a broader range of reasoning tasks.
@@ Appendix A Experimental Details
These estimations are performed for every problem in both the validation and test sets, providing the problem-specific parameters needed for our scaling plateau predictions.
@@ B.2 Detailed Derivation of the Inequality for N
We start from the condition for the scaling plateau (Eq ( 10)):
@@
For this inequality to be meaningful in the context of finding a positive $N$ where performance gain diminishes, we require the right-hand side to be less than 1 (since $(1-p_{x})^{N}$ will be less than 1 for $N\geq 1$ and $10<p_{x}<$ 1). If $\frac{\epsilon}{F_{\max}\cdot p_{x}}\geq 1$, it means that even for $N=0$ (or $N=1$, depending on interpretation), the marginal gain $F_{\max}\cdot p_{x}$ is already less than or equal to $\epsilon$, implying the plateau is reached immediately. Thus, we assume $10<\frac{\epsilon}{F_{\max}\cdot p_{x}}<$ 1, which implies $\epsilon<F_{\max}\cdot p_{x}$.
@@ Appendix C Additional Experimental Results
Figure LABEL:fig:acc_curve presents accuracy scaling curves across four benchmarks under different scaling strategies (parallel vs. sequential) and model sizes (1.5B vs. 7B). Across all settings, we observe a consistent trend: initial improvements in accuracy with increased generations per problem ($N$), followed by diminishing returns, forming a clear empirical scaling plateau.
In addition to the fact that the scaling plateau phenomenon is consistent and universal across tasks, models, and strategies, the above results further lead to several key insights: 1) Test-time scaling enables smaller models to rival or outperform larger, unscaled models. For example, the 1.5B model with sufficient scaling often approaches or exceeds the performance of the 7B vanilla baseline, particularly under parallel scaling. This highlights the practical value of test-time scaling as a lightweight alternative to model scale. 2) Parallel scaling consistently outperforms sequential scaling in terms of final accuracy. While both strategies benefit from increased $N$, parallel scaling achieves better asymptotic performance and exhibits more stable behavior across all benchmarks. These findings not only align with our theoretical predictions but also suggest that scaling compute at test time, especially via parallel strategies, can be a powerful tool for improving performance without increasing model size.